# Fighting Fire with Fire (F3): A Training-free and Efficient Visual Adversarial Example Purification Method in LVLMs

### v2

@@ Abstract.
Recent advances in large vision-language models (LVLMs) have showcased their remarkable capabilities across a wide range of multimodal vision-language tasks. However, these models remain vulnerable to visual adversarial attacks, which can substantially compromise their performance. Despite their potential impact, the development of effective methods for purifying such adversarial examples has received relatively limited attention. In this paper, we introduce F3, a novel adversarial purification framework that employs a counterintuitive “fighting fire with fire” strategy: intentionally introducing simple perturbations to adversarial examples to mitigate their harmful effects. Specifically, F3 leverages cross-modal attentions derived from randomly perturbed adversary examples as reference targets. By injecting noise into these adversarial examples, F3 effectively refines their attention, resulting in cleaner and more reliable model outputs. Remarkably, this seemingly paradoxical approach of employing noise to counteract adversarial attacks yields impressive purification results. Furthermore, F3 offers several distinct advantages: it is training-free and straightforward to implement, and exhibits significant computational efficiency improvements compared to existing purification methods. These attributes render F3 particularly suitable for large-scale industrial applications where both robust performance and operational efficiency are critical priorities. The code will be made publicly available.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Intrusion/anomaly detection and malware mitigation
@@ 1. Introduction
(2) Our method leverages an innovative approach by introducing random perturbations to adversarial examples for estimating the direction of clean attention. TThis estimated direction serves as a reference guide for F3’s additive noise generation, yielding a method that is novel, computationally efficient, and training-free.
@@ 2.1. Large Vision-Language Models (LVLMs)
Large vision-language models are built on powerful vision encoders (Dosovitskiy et al., 2020; Fang et al., 2023) and large language models (Chiang et al., 2023; Chung et al., 2024; Touvron et al., 2023; Zhang et al., 2022a), with the two components integrated through a vision-language projector. The process involves extracting features from input images using the vision encoder, which are then encoded by the vision-language projector into tokens compatible with the large language model’s input space. These image tokens, combined with text tokens, are processed by the large language model to generate responses. Widely used vision-language projectors include query transformer (Q-Former) (Li et al., 2023; Dai et al., 2024; Zhu et al., 2023) and multi-layer perceptron (MLP) (Liu et al., 2024b; Shi et al., 2024; Wang et al., 2024; Bai et al., 2023, 2025).
@@ 2.2. Adversarial Attacks and Defense Mechanisms
Adversarial examples are generated by strategically introducing small, carefully crafted perturbations to inputs, causing models to produce erroneous outputs. Early studies on adversarial attacks focused primarily on unimodal vision models, employing methods such as FGSM (Goodfellow et al., 2014b), PGD (Madry et al., 2017), JSMA (Papernot et al., 2016), DeepFool (Moosavi-Dezfooli et al., 2016), and the Carlini & Wagner (C&W) attack (Carlini and Wagner, 2017). However, recent research has demonstrated that LVLMs are equally vulnerable to such attacks. For example, Attack-Bard (Dong et al., 2023) generates adversarial examples across multiple surrogate models, successfully targeting ChatGPT-4 and Bard. Similarly, Carlini et al. (Carlini et al., 2024) utilized visual adversarial examples to trick LVLMs into producing harmful statements. To mitigate these threats, various defense mechanisms have been explored. Adversarial examples can be partially neutralized through techniques such as random resizing and padding (Xie et al., 2018), image super-resolution (Mustafa et al., 2019), and image compression (Jia et al., 2019), among others (Kang et al., 2021; Ho and Vasconcelos, 2022). Some studies have focused on detecting adversarial examples without purification (Zhang et al., 2024), while other research has explored the use of generative models for adversarial example purification. For instance, PixelDefend (Song et al., 2018) leverages PixelCNN (Van den Oord et al., 2016), Defense-GAN employs GAN-based architectures (Samangouei et al., 2018; Goodfellow et al., 2014a), and DiffPure utilizes diffusion models (Nie et al., 2022; Ho et al., 2020). However, these methods often apply generic filters or rely on separate models for image processing, without considering the specific inference model used. This limitation can lead to suboptimal robustness and high computational costs associated with training purification models and performing LVLM inference. In contrast, our work bridges this critical gap by specifically investigating adversarial example purification within the context of LVLMs. We present a training-free, efficient, and model-agnostic approach that achieves this through deliberately introducing noise into adversarial examples during inference.
@@ 3. F3: Fighting Fire with Fire
(1) $\mathcal{L}_{\text{C\&W}}(x_{i},x_{i}^{\prime},x_{t})=\mathcal{L}_{\text{LVLM}}(x_{i}^{\prime},x_{t})-c\times||x_{i}-x_{i}^{\prime}||_{2}.$
@@ 3.1. Cross-Modal Attention in LVLMs
Building on insights from PIP (Zhang et al., 2024), we focus specifically on the attention patterns within the LLM during the generation of the first token. In decoder-only LLMs, when generating this initial token, the model calculates attention weights between the token and all preceding $M+N$ tokens, using these weights to aggregate information (Vaswani et al., 2017). Our analysis centers on the cross-modal attention between the first response token and the visual tokens $I$, which we define as $\mathbf{A}(x_{i},x_{t},f)$ (abbreviated as $\mathbf{A}(x_{i})$). This attention tensor typically has dimensions $(L,H,M)$, where $f$ represents the LVLM, $L$ is the number of layers in the LLM, and $H$ is the number of attention heads. Cross-modal attention plays a critical role in extracting visual information for multimodal tasks during response generation, making it particularly relevant to studies on visual adversarial attacks and purification.
@@ 3.2. Cross-Modal Attention Differs Between Clean and Adversarial Examples
We then investigated the impact of adversarial attacks on cross-modal attention $\mathbf{A}$, as defined in Sec. 3.1. Our analysis revealed significant disparities between clean and adversarial examples in terms of $\mathbf{A}$. To facilitate visual comparison, we applied a maxima operation across the multi-head attention dimensions to project $\mathbf{A}$ into a two-dimensional representation. As illustrated in Fig. 2, there are notable differences between the attention of clean examples ($\mathbf{A}(x_{i})$) and adversarial examples ($\mathbf{A}(x_{i}^{\prime})$). Specifically, we observed an increase in attention values for the 8th token following the attacks. Similar patterns were consistently observed across various attack methods (e.g., PGD, as shown in Fig. 3) and different question types (e.g., yes/no questions in Fig. 4 and number-related questions in Fig. 5), with additional details provided in Appendix C. The differences in cross-modal attentions between clean and adversarial examples are consistent and can be quantitatively assessed using metrics such as mean-square error (MSE) and Kullback-Leibler (KL) divergence, as presented in the cells where column “Adv” intersects with rows “MSE” and “KL” of Tab. 1. These findings suggest potential strategies for mitigating or purifying adversarial examples.
$\gamma$ $\sim\mathcal{U}[0,\gamma_{\infty}].$
@@ 3.6. F3-v3: Finer Control of Purifying Noise
(5) $g=\nabla_{x_{i}^{\prime}}$ $||\mathbf{A}(x_{i}^{\prime},x_{t},f)-\mathbf{A}(\mathcal{R}(x_{i}^{\prime},\alpha_{\infty}),x_{t},f)||_{2}.$
@@
As shown in Eq. 5, the gradient $g$ is calculated with respect to reference attention using randomly perturbed adversarial examples. Equation 6 normalizes this gradient within the range $[0,1]$. To address the significant variability in the normalized gradient values ($g_{\text{norm}}$), Eq. 7 further amplifies $g_{\text{norm}}$ by dividing it by its mean value $\overline{g_{\text{norm}}}$, where $g_{\text{max}}$ and $g_{\text{min}}$ represent the maximum and minimum gradient values, respectively. Subsequently, noise is applied at each pixel location based on both the direction and magnitude of the gradient.
@@ 5. Conclusion
Appendix A Limitation
@@
Appendix B More Details, Results and Analysis for F3
B.1. Detailed Experimental Setup for Adaptive Attack Evaluation
@@
B.2. Distribution of question types in subdatasets
@@
B.3. Are the Attacks We Use Strong Enough?
@@
B.4. Generalizability of F3 over Different Attack Methods and Configurations
@@
B.5. Using Different Functions to Calculate the Noise via Cross-Modal Attention
@@
Appendix C Additional Visualization Results of Images and Cross-Modal Attentions

### v3

@@ Abstract.
Recent advances in large vision-language models (LVLMs) have showcased their remarkable capabilities across a wide range of multimodal vision-language tasks. However, these models remain vulnerable to visual adversarial attacks, which can substantially compromise their performance. In this paper, we introduce F3, a novel adversarial purification framework that employs a counterintuitive “fighting fire with fire” strategy: intentionally introducing simple perturbations to adversarial examples to mitigate their harmful effects. Specifically, F3 leverages cross-modal attentions derived from randomly perturbed adversary examples as reference targets. By injecting noise into these adversarial examples, F3 effectively refines their attention, resulting in cleaner and more reliable model outputs. Remarkably, this seemingly paradoxical approach of employing noise to counteract adversarial attacks yields impressive purification results. Furthermore, F3 offers several distinct advantages: it is training-free and straightforward to implement, and exhibits significant computational efficiency improvements compared to existing purification methods. These attributes render F3 particularly suitable for large-scale industrial applications where both robust performance and operational efficiency are critical priorities. The code is available at https://github.com/btzyd/F3.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758152isbn: 979-8-4007-2035-2/2025/10ccs: Security and privacy Intrusion/anomaly detection and malware mitigation
@@ 1. Introduction
(2) Our method leverages an innovative approach by introducing random perturbations to adversarial examples for estimating the direction of clean attention. This estimated direction serves as a reference guide for F3’s additive noise generation, yielding a method that is novel, computationally efficient, and training-free.
@@ 2.1. Large Vision-Language Models (LVLMs)
Large vision-language models are built on powerful vision encoders (Dosovitskiy et al., 2020; Fang et al., 2023; Zhang et al., 2025a) and large language models (Chiang et al., 2023; Chung et al., 2024; Touvron et al., 2023; Zhang et al., 2022a), with the two components integrated through a vision-language projector. The process involves extracting features from input images using the vision encoder, which are then encoded by the vision-language projector (Zhu et al., 2025; Zhang et al., 2025c) into tokens compatible with the large language model’s input space. These image tokens, combined with text tokens, are processed by the large language model to generate responses. Widely used vision-language projectors include Q-Former (Li et al., 2023; Dai et al., 2024; Zhu et al., 2023) and multi-layer perceptron (MLP) (Liu et al., 2024b; Shi et al., 2024; Wang et al., 2024; Bai et al., 2023, 2025).
@@ 2.2. Adversarial Attacks and Defense Mechanisms
Adversarial examples are generated by strategically introducing small, carefully crafted perturbations to inputs, causing models to produce erroneous outputs. Early studies on adversarial attacks focused primarily on unimodal vision models, employing methods such as FGSM (Goodfellow et al., 2014b), PGD (Madry et al., 2017), JSMA (Papernot et al., 2016), DeepFool (Moosavi-Dezfooli et al., 2016), and the Carlini & Wagner (C&W) attack (Carlini and Wagner, 2017). However, recent research has demonstrated that LVLMs are equally vulnerable to such attacks. For example, Attack-Bard (Dong et al., 2023) generates adversarial examples across multiple surrogate models, successfully targeting ChatGPT-4 and Bard. Similarly, Carlini et al. (Carlini et al., 2024) utilized visual adversarial examples to trick LVLMs into producing harmful statements. In addition, QAVA (Zhang et al., 2025b) is a query-agnostic visual attack method. To mitigate these threats, various defense mechanisms have been explored. Adversarial examples can be partially neutralized through techniques such as random resizing and padding (Xie et al., 2018), image super-resolution (Mustafa et al., 2019), and image compression (Jia et al., 2019), among others (Kang et al., 2021; Ho and Vasconcelos, 2022). Some studies have focused on detecting adversarial examples without purification (Zhang et al., 2024a), while other research has explored the use of generative models for adversarial example purification. For instance, PixelDefend (Song et al., 2018) leverages PixelCNN (Van den Oord et al., 2016), Defense-GAN employs GAN-based architectures (Samangouei et al., 2018; Goodfellow et al., 2014a), and DiffPure utilizes diffusion models (Nie et al., 2022; Ho et al., 2020). However, these methods often apply generic filters or rely on separate models for image processing, without considering the specific inference model used. This limitation can lead to suboptimal robustness and high computational costs associated with training purification models and performing LVLM inference. In contrast, our work bridges this critical gap by specifically investigating adversarial purification within the context of LVLMs. We present a training-free, efficient, and model-agnostic approach that achieves this through deliberately introducing noise into adversarial examples during inference.
@@ 3. F3: Fighting Fire with Fire
[…]
@@ 3.1. Cross-Modal Attention in LVLMs
Building on insights from PIP (Zhang et al., 2024a) and DHCP (Zhang et al., 2024b), we focus specifically on the attention patterns within the LLM during the generation of the first token. In decoder-only LLMs, when generating this initial token, the model calculates attention weights between the token and all preceding $M+N$ tokens, using these weights to aggregate information (Vaswani et al., 2017). Our analysis centers on the cross-modal attention between the first response token and the visual tokens $I$, which we define as $\mathbf{A}(x_{i},x_{t},f)$ (abbreviated as $\mathbf{A}(x_{i})$). This attention tensor typically has dimensions $(L,H,M)$, where $f$ represents the LVLM, $L$ is the number of layers in the LLM, and $H$ is the number of attention heads. Cross-modal attention plays a critical role in extracting visual information for multimodal tasks during response generation, making it particularly relevant to studies on visual adversarial attacks and purification.
@@ 3.2. Cross-Modal Attention Differs Between Clean and Adversarial Examples
We then investigated the impact of adversarial attacks on cross-modal attention $\mathbf{A}$, as defined in Sec. 3.1. Our analysis revealed significant disparities between clean and adversarial examples in terms of $\mathbf{A}$. To facilitate visual comparison, we applied a maxima operation across the multi-head attention dimensions to project $\mathbf{A}$ into a two-dimensional representation. As illustrated in Fig. 2, there are notable differences between the attention of clean examples ($\mathbf{A}(x_{i})$) and adversarial examples ($\mathbf{A}(x_{i}^{\prime})$). Specifically, we observed an increase in attention values for the 8th token following the attacks. Similar patterns were consistently observed across various attack methods and different question types and number-related questions, with additional details provided in Appendix. The differences in cross-modal attentions between clean and adversarial examples are consistent and can be quantitatively assessed using metrics such as mean-square error (MSE) and Kullback-Leibler (KL) divergence, as presented in the cells where column “Adv” intersects with rows “MSE” and “KL” of Tab. 1. These findings suggest potential strategies for mitigating or purifying adversarial examples.
@@ 3.6. F3-v3: Finer Control of Purifying Noise
[…]
@@
As shown in Eq. 5, the gradient $g$ is calculated with respect to reference attention using randomly perturbed adversarial examples. Equation 6 normalizes this gradient within the range $[0,1]$. To address the significant variability in the normalized gradient values ($g_{\text{norm}}$), Eq. 7 further amplifies $g_{\text{norm}}$ by dividing it by its mean value $\text{avg}({g_{\text{norm}}})$, where $g_{\text{max}}$ and $g_{\text{min}}$ represent the maximum and minimum gradient values, respectively. Subsequently, noise is applied at each pixel location based on both the direction and magnitude of the gradient.
@@ 5. Conclusion
Acknowledgements.
This work was supported by Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the National Key R&D Program of China (2023YFB4502200), the National Natural Science Foundation of China (No. 62376024, 62325405, 62104128, 62203257, 62031017, 62406159, U21B2031), Tsinghua University Initiative Scientific Research Program, Beijing National Research Center for Information Science, Technology (No. BNR2024TD03001) and Beijing Innovation Center for Future Chips.
Appendix A Experimental setup for attack and defense
Appendix B Limitation
@@
Appendix C More Details, Results and Analysis for F3
C.1. Detailed Experimental Setup for Adaptive Attack Evaluation
@@
C.2. Distribution of question types in subdatasets
@@
C.3. Are the Attacks We Use Strong Enough?
@@
C.4. Generalizability of F3 over Different Attack Methods and Configurations
@@
C.5. Using Different Functions to Calculate the Noise via Cross-Modal Attention
@@
Appendix D Additional Visualization Results of Images and Cross-Modal Attentions